### v2

# LLM-Assisted Relevance Assessments: When Should We Ask LLMs for Help?

Abstract.
Test collections are information retrieval tools that allow researchers to quickly and easily evaluate ranking algorithms. While test collections have become an integral part of IR research, the process of data creation involves significant effort in manual annotations, which often makes it very expensive and time-consuming. Thus, test collections could become too small when the budget is limited, which may lead to unstable evaluations. As a cheaper alternative, recent studies have proposed the use of large language models (LLMs) to completely replace human assessors. However, while LLMs may seem to somewhat correlate with human judgments, their predictions are not perfect and often show bias. Thus a complete replacement with LLMs is argued to be too risky and not fully reliable.
Thus, in this paper, we propose LLM-Assisted Relevance Assessments (LARA), an effective method to balance manual annotations with LLM annotations, which helps to build a rich and reliable test collection even under a low budget. We use the LLM’s predicted relevance probabilities to select the most profitable documents to manually annotate under a budget constraint. With theoretical reasoning, LARA effectively guides the human annotation process by actively learning to calibrate the LLM’s predicted relevance probabilities. Then, using the calibration model learned from the limited manual annotations, LARA debiases the LLM predictions to annotate the remaining non-assessed data. Empirical evaluations on TREC-7 Ad Hoc, TREC-8 Ad Hoc, TREC Robust 2004, and TREC-COVID datasets show that LARA outperforms alternative solutions under almost any budget constraint.
offline evaluation; relevance judgments; test collections
ccs: Information systems Test collections
1. Introduction
Test collections are laboratory tools that allow Information Retrieval (IR) systems to be reliably evaluated. A collection consists of relevance labels, which are annotations of whether a search result (we refer to as a document) is relevant to the searcher’s question (we refer to as a topic). Labels may come from various sources, but the most accurate option is to use the gold assessor labels (Bailey et al., 2008). In test collections, these labels originate from relevance assessors who develop their own topic and annotate whether documents are relevant to the topic, thereby establishing the ground truth. The relevance labels are then used to evaluate ranking algorithms easily (Thakur et al., 2021).
The Text REtrieval Conference (TREC) exemplifies community efforts to build large test collections with manual (gold) labels (Voorhees, 2005; Roberts et al., 2021; Craswell et al., 2024; Upadhyay et al., 2024b). Although these manual annotations are reliable, a major problem with this process is that they are costly and time-consuming. When the annotation budget is limited, test collections could become too small, potentially leading to unstable evaluation results. Over the years, various document selection techniques have been proposed to use human effort more efficiently (Cormack et al., 1998; Aslam et al., 2006; Aslam and Yilmaz, 2007; Li and Kanoulas, 2017), but the problem of data size has always persisted.
To solve the problem of annotation cost, some recent studies have proposed to use Large Language Models (LLMs) to replace the annotation process (Faggioli et al., 2023; Thomas et al., 2024; Qin et al., 2024). LLM-based assessments can drastically increase the size of test collections since the budget is no longer a problem. Moreover, studies have shown that LLM annotations perform almost always better than other low-cost alternatives like crowd-sourcing (Thomas et al., 2024). However, while these studies have shown some success, we can infer from their results that LLMs do not perfectly match the ground truth judgments. Particularly, LLM judgments are often biased (i.e., either too liberal or too strict (Alaofi et al., 2024; Thomas et al., 2024)). Even if a well-performing LLM or prompt is found on one dataset, there is no guarantee that it will perform the same in practice, due to differences in tasks, searchers, topics, or documents. Relying entirely on LLMs may also lead to theoretical challenges like the risk of overfitting to LLM-based metrics (Clarke and Dietz, 2024). As a result, relying solely on LLMs to build test collections is recently argued to be risky (Soboroff, 2024; Clarke and Dietz, 2024). This tension creates a dilemma: manual judgments are ground truth and high-quality but expensive; LLM judgments are cheap and scalable but error-prone and unreliable.
To address the dilemma between LLMs and humans, our work proposes LLM-Assisted Relevance Assessments (LARA) to achieve a reliable, yet budget-friendly annotation procedure. Given a limited budget, LARA uses LLMs to choose the most effective documents for manual annotation. Specifically, LARA uses LLMs’ output token probabilities to find their most uncertain predictions that should be manually annotated. Then, the rest of the documents (easy and more confident ones) can be judged by LLMs. While directly using LLM relevance probability as uncertainty is an effective method to find documents for manual annotation, we argue and show that these LLM predictions are often biased and are not optimal. Therefore, we propose an algorithm that actively learns to calibrate the LLM relevance probabilities, while continuously identifying the most effective documents to annotate manually. The remaining documents are then annotated using the calibrated LLM predictions, adjusted by the calibration model learned from human assessments. Thus, in our algorithm, calibrated LLM predictions guide manual annotations, and the gathered manual annotations in turn refine LLM predictions. Moreover, while most existing algorithms for efficient manual annotations are only suited for binary relevance labels (Cormack et al., 1998; Aslam et al., 2006; Li and Kanoulas, 2017), our LARA algorithm can deal with the general graded relevance framework (e.g., relevance levels 0/1/2), where the binary relevance setting is only a special case of our framework. Finally, we conduct comprehensive experiments on TREC-7 Ad Hoc, TREC-8 Ad Hoc, TREC Robust 2004, and TREC COVID datasets, where we show that our LARA algorithm almost constantly outperforms the tested alternative methods under various budget constraints. We are also the first to study and compare the performance of LLM-based methods and manual methods under different budget constraints.
2. Related Works
2.1. Efficient Human Annotations
When a document collection is large, annotating all topic-document pairs available is difficult due to the immense human labor required. To solve this problem, various methods have been proposed to select subsets of documents for human annotation (Cormack et al., 1998; Aslam et al., 2006; Aslam and Yilmaz, 2007; Li and Kanoulas, 2017; Voorhees, 2018).
Depth-k pooling is a common method for building test collections by judging only the union of the top-k retrieved documents from each system. Move-to-Front (MTF)(Cormack et al., 1998) adaptively picks unjudged documents and updates system priorities based on relevance, while multi-armed bandits (MAB)(Losada et al., 2016) alternates between the best system at each stage and random sampling. While these methods only rely on human annotations, thus providing reliable ground truth judgments, it is costly and time-consuming to build a large test collection. This often results in a small test set and inconsistent evaluations when there is not enough budget.
Active learning methods like Simple Active Learning (SAL)(Lewis and Gale, 1994) and Continuous Active Learning (CAL)(Cormack and Mojdeh, 2009), commonly used in e-Discovery (Cormack and Grossman, 2014; Oard et al., 2018; Cormack and Grossman, 2016; Stevenson and Bin-Hezam, 2023), have also been applied to building test collections (Rahman et al., 2020). SAL selects documents based on uncertainty sampling, while CAL focuses on likely-relevant documents. Rahman et al. (Rahman et al., 2020) also demonstrate hybrid approaches that use the trained model to judge unannotated documents. Although these hybrid methods share similar concepts with our study, our method employs a calibration model that leverages LLM predictions to guide annotation, achieving greater efficiency and effectiveness, as we show in our experiments.
2.2. Automated Relevance Assessments
Due to the high cost and effort needed for building test collections with ground truth labels, some have previously considered using cheaper third-party assessors (Alonso and Baeza-Yates, 2011; Eickhoff et al., 2012). However, working with these third-party labelers, especially crowd workers, can often suffer from misunderstanding the searcher needs, mistakes, collusion, “spammy” workers, and other biases (Clough et al., 2013; Thomas et al., 2022).
Researchers have also investigated fully automatic methods to create test collections (Dietz et al., 2022; Dietz and Dalton, 2020). Particularly, an increasing number of researchers are exploring the ability of LLMs in relevance judgments (Faggioli et al., 2023; Thomas et al., 2024; MacAvaney and Soldaini, 2023; Upadhyay et al., 2024a; Rahmani et al., 2024; Abbasiantaeb et al., 2024; Mehrdad et al., 2024; Alaofi et al., 2024; Upadhyay et al., 2024b).
These studies have shown the capability of LLMs in relevance assessments and have found that LLMs somewhat correlate with the ground truth labels and often perform better than crowd workers (Thomas et al., 2024). However, the relevance predictions by LLMs are often biased. For example, Alaofi et al. (Alaofi et al., 2024) report that LLMs tend to over-generate relevance labels compared to the ground truth. If we were to build a test collection only using LLMs, we would not be able to avoid such bias, making a full replacement risky (Soboroff, 2024; Clarke and Dietz, 2024). Even if a well-performing prompt is found on a certain dataset, there is no guarantee that it will work on an unseen task or data that we want to use for a test collection. To minimize such risks posed by LLMs, our LARA algorithm effectively uses partial ground truth judgments collected under a limited budget to actively calibrate the LLM predictions, building a more reliable dataset.
Upadhyay et al. (Upadhyay et al., 2024a) and Abbasiantaeb et al. (Abbasiantaeb et al., 2024) report the effect of using LLMs to fill relevance judgment holes that humans could not judge. Unlike these post-hoc procedures, our online approach leverages LLM predictions to guide which documents to manually judge, continually refining both the LLM calibration and our selection strategy.
3. LLM-Assisted Relevance Assessments
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 1. Showing the ground truth relevance probability ($p(y=1|\pi^{1})$) at each LLM predicted relevance probability ($\pi^{1}$) on TREC-7 Ad Hoc (top left), TREC-8 Ad Hoc (top right), TREC Robust04 (bottom left), and TREC-COVID (bottom right). The green line shows the correlation line if the correlation is perfect. For reference, the gray bars (secondary y-axis) show the number of predictions at each LLM probability value ($\pi^{1}$), showing where the LLM predictions are most concentrated.
3.1. Problem Setup
Let $D=\{x_{1},x_{2},...,x_{|D|}\}$ be the full dataset to annotate, where $x_{i}$ is the $i$-th data point (i.e., topic-document pair). We consider the graded relevance setting where each data point has an associated (ground-truth) human annotation of $y\in\{0,1,...,l\}$, representing a graded relevance label of $l+1$ levels. The binary relevance setting is a special case of our framework. We are given an annotation budget $B(\leq|D|)$, which is the maximum number of ground truth annotations that the human assessors can perform. The goal is to build a test collection that could accurately rank the evaluated IR systems in the order of performance.
While humans (topic developers) can provide the ground truth relevance labels, obtaining a sufficiently large fully human-annotated test collection can be infeasible due to the limited budget $B$. As a potential solution, recent studies suggest using LLM predictions to generate synthetic relevance assessments at scale (Faggioli et al., 2023; Thomas et al., 2024). However, while LLM annotations are cheap, they are not perfectly accurate and are unreliable (Alaofi et al., 2024; Thomas et al., 2024). This imbalance motivates our hybrid strategy: use human effort for the most effective annotations while relying on (calibrated) LLM assessments for the bulk of the data.
3.2. A Naive Uncertainty-Based Approach
Ideally, to minimize the number of annotation errors in the test collection, we want to spend our manual annotation budget on data points where LLMs are wrong. However, in practice, we cannot know this until the manual annotation has taken place. Thus, a straightforward strategy could be to focus manual effort on data points where the LLM is uncertain. Intuitively, if the LLM is unsure of a particular label, verifying that prediction would seem more effective than annotating data points where the LLM is confident.
To implement this idea, we propose to use the normalized LLM prediction probability $\pi^{k}=\pi(y=k|x)$ of each relevance level $k$. The probability $\pi^{k}$ is calculated for each data point $x_{i}$ as
(1) $\pi^{k}_{i}=\pi(y=k|x_{i})=\frac{p_{\text{LLM}}(k|x_{i})}{\sum_{j=0}^{l}p_{\text{LLM}}(j|x_{i})},$
where $p_{\text{LLM}}(k|x)$ is the probability of the LLM assigning the relevance level $k$ to the question “How relevant is the document to the topic?”. We express the LLM probability set for data point $x_{i}$ as $\pi_{i}=\{\pi_{i}^{0},\pi_{i}^{1},...,\pi_{i}^{l}\}$. The prompts are provided in Appendix A.
Given these probabilities, one naive approach is to identify the data points on which LLM is the most uncertain and annotate those manually, which we call the Naive method. For a graded relevance setting, a simple heuristic is to first determine the most likely predicted level $k^{\prime}$ and the second most likely predicted level $s^{\prime}$ for each data point:
(2) $k^{\prime}_{i}=\operatorname*{arg\,max}_{j}\pi_{i}^{j}\quad\text{and}\quad s^{\prime}_{i}=\operatorname*{arg\,max}_{j\neq k^{\prime}_{i}}\pi_{i}^{j}.$
Then, the LLM’s most uncertain data point $x_{i}$ would be where the margin $m^{\prime}_{i}=\pi_{i}^{k^{\prime}}-\pi^{s^{\prime}}_{i}$ is the smallest. Thus, the Naive method manually annotates $B$ data points with the smallest margins $m^{\prime}$, and the remaining data points are given the LLM predictions $\hat{y}=k^{\prime}$.
3.3. LLM-Assisted Relevance Assessments
Although it may seem like the Naive method would work well, LLMs’ uncertain data points are not necessarily the best data points to annotate, mainly because LLM predictions are often biased (Thomas et al., 2024). To show this and to develop a better method, we first introduce a theorem that shows the optimal inspection point for such a classification problem, built on the concept of uncertainty sampling.
Theorem 1 (Optimal Inspection Point).
Let $p(y=j|\pi)$ denote the true probability of relevance level $j$ given $\pi$. Let $k=\operatorname*{arg\,max}_{j}p(y=j|\pi)$ be the level with the highest probability and $s=\operatorname*{arg\,max}_{j\neq k}p(y=j|\pi)$ be the level with the second highest probability. Then, the greatest potential reduction in classification error is achieved by annotating the data point with the smallest margin
$m=p(y=k|\pi)-p(y=s|\pi),$
where we suppose that the remaining data are classified into $\hat{y}=\text{argmax}_{j}p(y=j|\pi)$. We share the proof upon acceptance. Appendix B proves this via the maximum uncertainty principle.
Theorem 1 indicates that the data point with the smallest naive margin $m^{\prime}$ may not be the best point to annotate unless $\pi$ is perfectly correlated with the true relevance probabilities $p(y=j|\pi^{j})$. To empirically show some examples of the correlation, Figure 1 shows the relationship between the Llama-3.1-70B-Instruct predictions $\pi^{1}$ and the true relevance probabilities $p(y=1|\pi^{1})$, on the TREC-7 Ad Hoc, TREC-8 Ad Hoc, TREC Robust04, and the TREC COVID dataset. Refer to Table 1 for statistics of each dataset. From the figure, we see that the correlation is far from perfect alignment, which shows that selecting documents based solely on the LLM’s raw probabilities may not pinpoint where the true uncertainty lies.
From Theorem 1, we know that we should instead annotate data points with the smallest true margin $m$. However, since the true probabilities $p(y=j|\pi)$ are unknown, we propose to learn this calibration mapping for each relevance level $j\in\{0,...,l\}$ while selecting reasonable and effective data points for manual annotation.
Now, we explain our LARA algorithm. A more detailed explanation is provided in Algorithm 1. First, the calibration is set to $\hat{p}(y|\pi)=\pi$. Thus, we start by manually annotating data points closest to $m^{\prime}=\pi_{i}^{k^{\prime}}-\pi^{s^{\prime}}_{i}$ like the Naive method. Then, using these initial manual annotations, we train a calibration estimator and gradually update the prediction $\hat{p}(y|\pi)$ for all relevance levels. The calibration can be performed by regressing the true relevance $y$ on the estimated relevance probability $\pi$. We then repeat the process of annotating data points with the smallest predicted margin
$\hat{m}=\hat{p}(y=k|\pi)-\hat{p}(y=s|\pi),$
where $k=\operatorname*{arg\,max}_{j}\hat{p}(y=j|\pi)$ and $s=\operatorname*{arg\,max}_{j\neq k}\hat{p}(y=j|\pi)$, and we iteratively update the calibration until budget $B$ is exhausted. Finally, for the data points that were not manually annotated, we assign $\hat{y}=\operatorname*{arg\,max}_{j}\hat{p}(y=j|\pi)$. This final step allows the data points that were not manually annotated to be given calibrated and debiased LLM predictions learned on limited ground-truth labels.
Algorithm 1 LARA: LLM-Assisted Relevance Assessments
1:Input: Dataset $D=\{x_{1},x_{2},...,x_{|D|}\}$, LLM predicted probability vectors $\{\pi_{i}\}_{i=1}^{|D|}$ where $\pi_{i}=\{\pi_{i}^{0},\pi_{i}^{1},\ldots,\pi_{i}^{l}\}$, budget $B$
2:Output: Truth labels $y_{A}$ and predicted annotations $\hat{y}_{D\setminus A}$
3:Initialize $\hat{p}(y=j|\pi)=\pi^{j}$ for all $j\in\{0,...,l\}$ and for all $x_{i}\in D$
4:$A\leftarrow\emptyset$
5:while $|A|<B$ do
6:     for $x_{i}\in D\setminus A$ do
7:         $k=\operatorname*{arg\,max}_{j}\hat{p}(y=j|\pi_{i})$
8:         $s=\operatorname*{arg\,max}_{j\neq k}\hat{p}(y=j|\pi_{i})$
9:         Compute the margin $\hat{m}_{i}=\hat{p}(y=k|\pi_{i})-\hat{p}(y=s|\pi_{i})$
10:     end for
11:     $x^{*}\leftarrow\arg\min_{x_{i}\in D\setminus A}\hat{m}_{i}$
12:     $y^{*}\leftarrow\text{ManualAnnotation}(x^{*})$
13:     $A\leftarrow A\cup\{(x^{*},y^{*})\}$
14:     UpdateCalibrator$(\hat{p},A)$
15:end while
16:$y_{A}\leftarrow\{y^{*}:(x^{*},y^{*})\in A\}$
17:for $x_{i}\in D\setminus A$ do
18:     $\hat{y}_{i}=\operatorname*{arg\,max}_{j}\hat{p}(y=j|\pi_{i})$
19:end for
20:return $(y_{A},\hat{y}_{D\setminus A})$
3.4. Accessibility for Real-World Implementation
While our LARA algorithm appears highly effective, it may lack practicality. This is because, in the proposed algorithm, LARA is not suited for having multiple (gold) assessors. In venues that build these test collections, coordinators usually assign multiple assessors to come up with topics and assess documents (Voorhees and Harman, 2000). Usually, each assessor comes up with a number of their own topics based on their own interests and then judges the relevance of the given documents. However, if the LARA algorithm is used by multiple assessors, each assessor would have to assess documents whenever their topic is chosen; this requires all assessors to be working at the same time.
To solve this problem, we bundle the topics into groups, based on the number of assessors $n$. Then, we divide the budget $B$ by the number of assessors; each group of topics would be assigned with the budget $\frac{B}{n}$. We only sample from one group until the budget $\frac{B}{n}$ is exhausted, then we move on to the next group. This way, assessors do not have to be working at the same time. In the experiments, we denote the number of assessors as $\text{LARA}(n)$. For example, we write $\mathrm{LARA}(n=3)$ if there are 3 annotators. When each topic has a different annotator, we write $\mathrm{LARA}(n=N)$ where $N$ is the total number of topics. We test the grouped methods in our experiments to see how the performance of LARA changes with different numbers of annotators.
4. Experiment Setup
In this section, we conduct experiments to see the performance of LARA compared to the alternative methods of building test collections under different budget constraints $B$. In the experiments, the budget is shown as a ratio against the total number of assessments in the full test collection (i.e., $B/|D|$). The statistics of the datasets are provided in Table 1. All datasets are originally constructed via the Depth-$k$ method. Each dataset contains runs submitted by the participating systems, where each run gives a ranking of documents given a topic. We aim to build a test collection that can accurately rank these systems in the order of performance.
Table 1. Statistics of the datasets used in our experiments.
Dataset Relevance Levels #Data Points #Systems
TREC-7 Ad Hoc 0, 1 80,346 103
TREC-8 Ad Hoc 0, 1 86,829 131
TREC Robust04 0, 1, 2 311,392 110
TREC-COVID 0, 1, 2 60,075 126
4.1. Evaluation Metrics
For each system performance evaluation, we use the Mean Average Precision (MAP) score for the binary relevance experiments and the Normalized Discounted Cumulative Gain (NDCG) score for the graded relevance experiments. Then, based on the MAP/NDCG scores, we rank the participating systems. We consider the system rankings evaluated by the full test collection to be the ground truth ranking. We evaluate the system rankings created under budget by computing Kendall’s $\tau$, which will be our main metric to evaluate the method’s performance. If the evaluated ranking is identical to the ground truth ranking built by the full test collection, the score will be 1, which would mean that it is a perfect ranking.
In addition to the main performance evaluation through Kendall’s $\tau$, we report the Maximum Drop, which shows the maximum decrease in rank for a system in the ground truth ranking compared to the evaluated ranking. The Maximum Drop has been reported in recent studies because Kendall’s $\tau$ scores over large lists can indicate general concordance despite a few items with large rank differences (Voorhees, 2018; Voorhees et al., 2022). Higher Maximum Drop values indicate that at least one system is being unfairly treated in the ranking.
One crucial advantage of our algorithm, beyond selecting effective documents to manually annotate, is that they can minimize the errors of the LLM annotations. Specifically, LARA calibrates the LLM predictions by learning a model $\hat{p}(y|\pi)$ from the collected human annotations. Moreover, in LARA (and in the Naive method), we try to give ground-truth labels to the most uncertain data points, which also helps minimize the LLM prediction errors. Thus, as exploratory analyses, we show the performance increase of LLM annotations with LARA by evaluating via overlap score (Voorhees, 1998).
(3) Overlap=# of True Positives# of True Positives+# of False $\text{Overlap}=\frac{\text{\# of True Positives}}{\text{\# of True Positives}+\text{\# of False Predictions}}\,\,\,\,\,,$
where True Positives are when ground truth and LLM both judge the same relevance and are not non-relevant (i.e., $y_{i}=\hat{y}_{i}\geq 1$), and the False Predictions are all the false predictions (i.e., $y_{i}\neq\hat{y}_{i}$). We use the overlap metric over Cohen’s Kappa because the number of non-relevant labels greatly exceeds the number of relevant documents in the experimented test collections.
4.2. Comparison Methods
We compare the LARA algorithm with several baseline approaches. We first introduce methods from previous studies that only use manual assessments. Note that since most of these existing manual methods are only suited for binary relevance levels, we binarize the relevance levels in the learning procedure by setting a threshold at $\frac{l}{2}$, but we keep the graded relevance levels for final annotation.
Depth-k Pooling. Depth-k Pooling is a method with only human assessments. It assesses the top-k documents at each rank for each topic for all systems until the budget is empty.
Move-to-Front Pooling (MTF) (Cormack et al., 1998). In the MTF algorithm, all systems start with equal priority. On each round, it selects the first unjudged document provided by the highest-priority system to judge. If relevant, it continues; if not, the system’s priority is lowered, and the next system is selected.
Multi-armed Bandits (MAB) (Losada et al., 2016). This method treats selection as a multi-armed bandit problem, randomly choosing between the top-priority system or sampling from the entire collection. For the MAB baseline, we used the best-performing method, MaxMean Non-Stationary (MM-NS), with default settings from (Losada et al., 2016).
Continuous Active Learning (CAL) (Cormack and Mojdeh, 2009). CAL selects the most likely relevant data point to annotate and continuously trains a regression model. We compare CAL(human), which consists of only manual annotations, and CAL(hybrid), which also predicts the remaining data points with the trained regression model.
Simple Active Learning (SAL) (Lewis and Gale, 1994). Similar to CAL, SAL iteratively trains a regression model, but SAL instead selects data points based on uncertainty sampling. We compare both SAL(human) and SAL(hybrid) in the experiments.
In addition to the above manual methods, we add three more baselines that incorporate LLM predictions:
LLM-Only. This method uses only LLM judgments to build a test collection.
Random. This method uses both LLM annotations and human annotations. $B$ randomly chosen data points are manually annotated and the rest are judged by LLMs. This method can also be seen as using LLMs to fill the holes of manual judgments (Upadhyay et al., 2024a; Abbasiantaeb et al., 2024).
Naive. This method uses both humans and LLMs to build a test collection. It chooses the LLM’s most uncertain data points to manually annotate by naively using LLM relevance probabilities $\pi$, then annotates the rest using LLM predictions. Further details of this method are provided in Section 3.1. Note that the Naive method was never introduced by any of the previous studies; we compare this method as a baseline only to visualize the advantage of our LARA calibration algorithm.
We investigate the effect of having different numbers of annotators. Thus, we show results of LARA($\boldsymbol{n=1}$), LARA($\boldsymbol{n=3}$), and LARA($\boldsymbol{n=N}$). More on the number of annotators is in Section 3.4. Logistic regression is used as the calibration model for LARA.
The base prompt used in the experiments is provided in Appendix A. To study the effect of the prompt on LARA, we also share the results with the rationale prompt based on Upadhyay et al. (Upadhyay et al., 2024b) and the utility prompt based on Thomas et al. (Thomas et al., 2024) as an extensive study. The prompts are slightly modified from the original to better align with our task [^1] We share the full rationale and utility prompts upon acceptance.. Methods that use these prompts are named in the format MethodName(PromptType), like LARA(rationale) and LARA(utility), where the LARA methods here use LARA($n=N$) for the number of annotators.
Unless specified otherwise, all LLM-based baseline methods use Llama-3.1-70B-Instruct [^2] https://ai.meta.com/blog/meta-llama-3-1/. Temperature is set to $3.0$, and top-$k$ is set to 50, so the LLM can output the probabilities of answering various relevance grades. To study the effect of LLM performance on LARA, we also show results using the lightweight variant of the base LLM, Llama-3.1-8B-Instruct, as an ablation study. Methods that use the 8B-version are named LLM-Only (Llama-8B), Random (Llama-8B), Naive (Llama-8B), and LARA (Llama-8B), where LARA (Llama-8B) uses LARA($n=N$) for the number of annotators.
Table 2. Kendall’s Tau correlation scores (larger is better) for the ranked systems based on MAP scores across different methods and different ratios of manual annotation budget. The first blocks show the main results, where for each column, the top three highest scores are shown in bold, the fourth is in italics, and the fifth is underlined. For the methods based on the rationale and utility prompt, only the top score of each prompt is shown in bold. For the methods based on Llama-8B, the top score is highlighted in bold. Values in parentheses show the Maximum Drop in the system rank (smaller is better).
TREC-7 Ad Hoc
Ratio of Manual Annotation Budget ($B/|D|$)
Method $\frac{1}{512}$ $\frac{1}{256}$ $\frac{1}{128}$ $\frac{1}{64}$ $\frac{1}{32}$ $\frac{1}{16}$ $\frac{1}{8}$ $\frac{1}{4}$ $\frac{1}{2}$ all
Depth-k 0.495 (65) 0.624 (83) 0.704 (72) 0.751 (47) 0.838 (39) 0.879 (18) 0.922 (13) 0.931 (5) 0.973 (3) 1.000 (0)
MTF 0.508 (58) 0.578 (56) 0.657 (56) 0.694 (56) 0.675 (45) 0.764 (39) 0.810 (28) 0.837 (24) 0.920 (15) 1.000 (0)
MM-NS 0.405 (56) 0.405 (56) 0.465 (61) 0.747 (42) 0.802 (32) 0.813 (25) 0.874 (18) 0.923 (8) 0.961 (4) 1.000 (0)
CAL(human) 0.154 (59) 0.225 (65) 0.212 (74) 0.284 (73) 0.674 (63) 0.650 (39) 0.822 (25) 0.839 (33) 0.897 (17) 1.000 (0)
CAL(hybrid) 0.459 (71) 0.414 (74) 0.372 (71) 0.509 (85) 0.583 (50) 0.693 (26) 0.788 (27) 0.792 (52) 0.895 (21) 1.000 (0)
SAL(human) 0.471 (63) 0.493 (33) 0.460 (62) 0.534 (61) 0.660 (84) 0.784 (29) 0.786 (29) 0.830 (23) 0.887 (15) 1.000 (0)
SAL(hybrid) 0.317 (60) 0.467 (60) 0.342 (84) 0.347 (78) 0.324 (38) 0.784 (38) 0.827 (24) 0.843 (28) 0.899 (14) 1.000 (0)
LLM-Only 0.861 (42) 0.861 (42) 0.861 (42) 0.861 (42) 0.861 (42) 0.861 (42) 0.861 (42) 0.861 (42) 0.861 (42) 0.861 (42)
Random 0.862 (42) 0.861 (42) 0.861 (42) 0.861 (42) 0.864 (38) 0.865 (38) 0.873 (37) 0.885 (34) 0.911 (24) 1.000 (0)
Naive 0.861 (42) 0.862 (41) 0.863 (38) 0.864 (38) 0.868 (37) 0.883 (36) 0.905 (30) 0.923 (23) 0.938 (15) 1.000 (0)
LARA($n=1$) 0.869 (36) 0.877 (36) 0.894 (31) 0.900 (30) 0.910 (27) 0.917 (23) 0.930 (20) 0.946 (10) 0.973 (6) 1.000 (0)
LARA($n=3$) 0.875 (36) 0.882 (36) 0.894 (31) 0.900 (30) 0.907 (27) 0.916 (23) 0.929 (20) 0.942 (12) 0.980 (4) 1.000 (0)
LARA($n=N$) 0.875 (36) 0.884 (36) 0.893 (32) 0.903 (29) 0.912 (28) 0.922 (23) 0.936 (19) 0.941 (11) 0.966 (7) 1.000 (0)
LLM-Only(rationale) 0.837 (42) 0.837 (42) 0.837 (42) 0.837 (42) 0.837 (42) 0.837 (42) 0.837 (42) 0.837 (42) 0.837 (42) 0.837 (42)
Random(rationale) 0.837 (42) 0.837 (42) 0.840 (42) 0.839 (42) 0.842 (41) 0.842 (41) 0.852 (38) 0.865 (35) 0.906 (24) 1.000 (0)
Naive(rationale) 0.839 (42) 0.840 (42) 0.842 (41) 0.845 (41) 0.858 (39) 0.872 (36) 0.899 (30) 0.929 (22) 0.950 (16) 1.000 (0)
LARA(rationale) 0.846 (41) 0.857 (36) 0.866 (35) 0.870 (34) 0.881 (31) 0.915 (21) 0.928 (21) 0.942 (15) 0.982 (5) 1.000 (0)
LLM-Only(utility) 0.846 (40) 0.846 (40) 0.846 (40) 0.846 (40) 0.846 (40) 0.846 (40) 0.846 (40) 0.846 (40) 0.846 (40) 0.846 (40)
Random(utility) 0.846 (40) 0.846 (40) 0.847 (40) 0.847 (40) 0.848 (39) 0.847 (37) 0.857 (36) 0.873 (33) 0.912 (22) 1.000 (0)
Naive(utility) 0.845 (40) 0.845 (40) 0.847 (39) 0.853 (39) 0.866 (37) 0.881 (31) 0.905 (20) 0.926 (19) 0.948 (13) 1.000 (0)
LARA(utility) 0.863 (35) 0.873 (33) 0.880 (30) 0.891 (27) 0.892 (24) 0.908 (20) 0.926 (17) 0.940 (12) 0.971 (7) 1.000 (0)
LLM-Only (Llama-8B) 0.578 (93) 0.578 (93) 0.578 (93) 0.578 (93) 0.578 (93) 0.578 (93) 0.578 (93) 0.578 (93) 0.578 (93) 0.578 (93)
Random (Llama-8B) 0.579 (93) 0.577 (93) 0.579 (93) 0.579 (93) 0.580 (93) 0.581 (93) 0.595 (92) 0.621 (88) 0.696 (77) 1.000 (0)
Naive (Llama-8B) 0.578 (93) 0.579 (93) 0.579 (93) 0.581 (93) 0.585 (92) 0.595 (92) 0.605 (92) 0.643 (85) 0.730 (62) 1.000 (0)
LARA (Llama-8B) 0.643 (85) 0.676 (82) 0.715 (72) 0.737 (61) 0.746 (58) 0.756 (52) 0.790 (41) 0.859 (32) 0.924 (15) 1.000 (0)
TREC-8 Ad Hoc
Ratio of Manual Annotation Budget ($B/|D|$)
Method $\frac{1}{512}$ $\frac{1}{256}$ $\frac{1}{128}$ $\frac{1}{64}$ $\frac{1}{32}$ $\frac{1}{16}$ $\frac{1}{8}$ $\frac{1}{4}$ $\frac{1}{2}$ all
Depth-k 0.547 (95) 0.597 (74) 0.657 (99) 0.699 (88) 0.790 (62) 0.882 (22) 0.913 (19) 0.940 (12) 0.982 (6) 1.000 (0)
MTF 0.291 (90) 0.281 (90) 0.320 (97) 0.340 (99) 0.323 (97) 0.590 (93) 0.758 (43) 0.830 (28) 0.915 (18) 1.000 (0)
MM-NS 0.319 (98) 0.314 (100) 0.478 (76) 0.525 (92) 0.726 (42) 0.756 (45) 0.844 (43) 0.888 (23) 0.957 (13) 1.000 (0)
CAL(human) 0.459 (93) 0.560 (65) 0.509 (112) 0.547 (65) 0.505 (60) 0.671 (54) 0.722 (36) 0.773 (37) 0.891 (28) 1.000 (0)
CAL(hybrid) 0.590 (81) 0.627 (107) 0.494 (96) 0.627 (43) 0.535 (45) 0.688 (68) 0.726 (54) 0.754 (47) 0.889 (29) 1.000 (0)
SAL(human) 0.325 (99) 0.328 (68) 0.620 (88) 0.522 (65) 0.574 (70) 0.677 (45) 0.644 (42) 0.847 (33) 0.890 (28) 1.000 (0)
SAL(hybrid) 0.278 (86) 0.455 (75) 0.352 (86) 0.515 (60) 0.692 (53) 0.730 (39) 0.778 (35) 0.809 (44) 0.891 (25) 1.000 (0)
LLM-Only 0.788 (99) 0.788 (99) 0.788 (99) 0.788 (99) 0.788 (99) 0.788 (99) 0.788 (99) 0.788 (99) 0.788 (99) 0.788 (99)
Random 0.788 (98) 0.789 (98) 0.789 (98) 0.789 (97) 0.792 (96) 0.797 (94) 0.806 (93) 0.832 (89) 0.898 (73) 1.000 (0)
Naive 0.789 (98) 0.790 (96) 0.791 (95) 0.794 (94) 0.802 (93) 0.818 (89) 0.847 (84) 0.883 (72) 0.912 (49) 1.000 (0)
LARA($n=1$) 0.834 (87) 0.848 (79) 0.860 (65) 0.870 (55) 0.875 (38) 0.897 (19) 0.917 (16) 0.939 (13) 0.970 (11) 1.000 (0)
LARA($n=3$) 0.843 (84) 0.850 (72) 0.864 (60) 0.868 (50) 0.874 (37) 0.896 (18) 0.921 (16) 0.939 (13) 0.970 (11) 1.000 (0)
LARA($n=N$) 0.845 (82) 0.856 (72) 0.875 (60) 0.882 (50) 0.889 (36) 0.904 (20) 0.925 (18) 0.950 (12) 0.977 (9) 1.000 (0)
LLM-Only(rationale) 0.843 (90) 0.843 (90) 0.843 (90) 0.843 (90) 0.843 (90) 0.843 (90) 0.843 (90) 0.843 (90) 0.843 (90) 0.843 (90)
Random(rationale) 0.843 (90) 0.845 (90) 0.845 (90) 0.845 (89) 0.849 (88) 0.854 (88) 0.860 (87) 0.878 (84) 0.922 (46) 1.000 (1)
Naive(rationale) 0.846 (89) 0.846 (88) 0.849 (88) 0.850 (86) 0.858 (86) 0.876 (74) 0.892 (41) 0.935 (15) 0.959 (11) 1.000 (1)
LARA(rationale) 0.847 (86) 0.860 (79) 0.871 (72) 0.881 (60) 0.899 (38) 0.908 (19) 0.938 (18) 0.957 (13) 0.986 (5) 1.000 (1)
LLM-Only(utility) 0.871 (75) 0.871 (75) 0.871 (75) 0.871 (75) 0.871 (75) 0.871 (75) 0.871 (75) 0.871 (75) 0.871 (75) 0.871 (75)
Random(utility) 0.871 (74) 0.872 (74) 0.874 (74) 0.873 (73) 0.879 (72) 0.883 (71) 0.890 (68) 0.907 (56) 0.927 (22) 1.000 (1)
Naive(utility) 0.871 (74) 0.877 (74) 0.878 (72) 0.884 (68) 0.896 (60) 0.911 (40) 0.933 (17) 0.949 (12) 0.968 (7) 1.000 (1)
LARA(utility) 0.877 (67) 0.882 (50) 0.886 (44) 0.903 (23) 0.919 (18) 0.930 (19) 0.937 (19) 0.962 (13) 0.988 (5) 1.000 (1)
LLM-Only (Llama-8B) 0.676 (99) 0.676 (99) 0.676 (99) 0.676 (99) 0.676 (99) 0.676 (99) 0.676 (99) 0.676 (99) 0.676 (99) 0.676 (99)
Random (Llama-8B) 0.676 (99) 0.676 (99) 0.677 (99) 0.680 (99) 0.683 (99) 0.687 (99) 0.697 (99) 0.717 (98) 0.784 (96) 1.000 (1)
Naive (Llama-8B) 0.675 (99) 0.676 (99) 0.677 (99) 0.680 (99) 0.684 (99) 0.691 (99) 0.707 (98) 0.723 (97) 0.788 (95) 1.000 (1)
LARA (Llama-8B) 0.723 (91) 0.754 (90) 0.773 (84) 0.786 (83) 0.812 (77) 0.845 (70) 0.871 (66) 0.890 (45) 0.926 (18) 1.000 (1)
Table 3. Kendall’s Tau correlation scores (larger is better) for the ranked systems based on NDCG scores across different methods and different ratios of manual annotation budget. The first blocks show the main results, where for each column, the top three highest scores are shown in bold, the fourth is in italics, and the fifth is underlined. For the methods based on the rationale and utility prompt, only the top score of each prompt is shown in bold. For the methods based on Llama-8B, the top score is highlighted in bold. Values in parentheses show the Maximum Drop in the system rank (smaller is better).
TREC Robust04
Ratio of Manual Annotation Budget ($B/|D|$)
Method $\frac{1}{512}$ $\frac{1}{256}$ $\frac{1}{128}$ $\frac{1}{64}$ $\frac{1}{32}$ $\frac{1}{16}$ $\frac{1}{8}$ $\frac{1}{4}$ $\frac{1}{2}$ all
Depth-k 0.662 (44) 0.656 (40) 0.699 (39) 0.756 (37) 0.793 (34) 0.846 (26) 0.898 (18) 0.939 (12) 0.977 (5) 1.000 (0)
MTF 0.420 (60) 0.442 (56) 0.501 (59) 0.535 (57) 0.608 (49) 0.611 (48) 0.673 (41) 0.727 (33) 0.951 (10) 1.000 (0)
MM-NS 0.490 (57) 0.490 (57) 0.490 (57) 0.577 (56) 0.645 (44) 0.622 (47) 0.678 (41) 0.730 (32) 0.950 (10) 1.000 (0)
CAL(human) 0.662 (52) 0.775 (31) 0.779 (30) 0.857 (26) 0.852 (22) 0.856 (34) 0.914 (21) 0.874 (21) 0.914 (13) 1.000 (0)
CAL(hybrid) 0.695 (69) 0.710 (54) 0.787 (24) 0.832 (25) 0.834 (28) 0.900 (23) 0.868 (37) 0.873 (28) 0.919 (12) 1.000 (0)
SAL(human) 0.376 (54) 0.665 (39) 0.765 (43) 0.799 (26) 0.819 (39) 0.860 (19) 0.925 (17) 0.892 (18) 0.918 (11) 1.000 (0)
SAL(hybrid) 0.724 (45) 0.663 (41) 0.638 (55) 0.770 (35) 0.799 (34) 0.820 (27) 0.888 (17) 0.914 (18) 0.931 (11) 1.000 (0)
LLM-Only 0.948 (22) 0.948 (22) 0.948 (22) 0.948 (22) 0.948 (22) 0.948 (22) 0.948 (22) 0.948 (22) 0.948 (22) 0.948 (22)
Random 0.948 (22) 0.948 (22) 0.949 (22) 0.948 (22) 0.947 (22) 0.946 (22) 0.948 (22) 0.948 (21) 0.955 (21) 1.000 (0)
Naive 0.948 (22) 0.949 (22) 0.950 (22) 0.951 (22) 0.950 (22) 0.959 (23) 0.965 (21) 0.971 (19) 0.993 (18) 1.000 (0)
LARA($n=1$) 0.949 (22) 0.949 (22) 0.950 (22) 0.951 (22) 0.953 (22) 0.960 (22) 0.968 (22) 0.972 (21) 0.994 (11) 1.000 (0)
LARA($n=3$) 0.949 (22) 0.949 (22) 0.950 (22) 0.951 (22) 0.953 (22) 0.960 (22) 0.968 (22) 0.972 (21) 0.994 (11) 1.000 (0)
LARA($n=N$) 0.949 (22) 0.949 (22) 0.950 (22) 0.951 (22) 0.957 (22) 0.960 (22) 0.968 (21) 0.976 (18) 0.993 (10) 1.000 (0)
LLM-Only(rationale) 0.945 (23) 0.945 (23) 0.945 (23) 0.945 (23) 0.945 (23) 0.945 (23) 0.945 (23) 0.945 (23) 0.945 (23) 0.945 (23)
Random(rationale) 0.945 (23) 0.945 (22) 0.945 (22) 0.946 (22) 0.951 (22) 0.954 (22) 0.962 (21) 0.970 (18) 0.988 (10) 1.000 (0)
Naive(rationale) 0.945 (23) 0.945 (23) 0.945 (23) 0.946 (22) 0.952 (22) 0.955 (22) 0.962 (21) 0.971 (18) 0.987 (10) 1.000 (0)
LARA(rationale) 0.945 (22) 0.945 (22) 0.946 (22) 0.947 (22) 0.953 (22) 0.956 (22) 0.964 (21) 0.972 (18) 0.989 (10) 1.000 (0)
LLM-Only(utility) 0.947 (21) 0.947 (21) 0.947 (21) 0.947 (21) 0.947 (21) 0.947 (21) 0.947 (21) 0.947 (21) 0.947 (21) 0.947 (21)
Random(utility) 0.946 (21) 0.945 (21) 0.947 (20) 0.948 (20) 0.953 (20) 0.957 (20) 0.964 (19) 0.978 (17) 0.993 (10) 1.000 (0)
Naive(utility) 0.947 (21) 0.947 (21) 0.946 (20) 0.947 (20) 0.954 (20) 0.956 (20) 0.963 (19) 0.979 (17) 0.994 (10) 1.000 (0)
LARA(utility) 0.947 (21) 0.947 (21) 0.948 (20) 0.949 (20) 0.955 (20) 0.958 (20) 0.966 (19) 0.980 (17) 0.995 (10) 1.000 (0)
LLM-Only (Llama-8B) 0.894 (33) 0.894 (33) 0.894 (33) 0.894 (33) 0.894 (33) 0.894 (33) 0.894 (33) 0.894 (33) 0.894 (33) 0.894 (33)
Random (Llama-8B) 0.894 (33) 0.894 (33) 0.894 (33) 0.894 (33) 0.894 (33) 0.894 (33) 0.896 (32) 0.898 (32) 0.905 (28) 1.000 (0)
Naive (Llama-8B) 0.894 (33) 0.894 (33) 0.894 (33) 0.897 (33) 0.899 (33) 0.901 (32) 0.905 (32) 0.906 (29) 0.919 (24) 1.000 (0)
LARA (Llama-8B) 0.896 (33) 0.896 (33) 0.896 (33) 0.900 (33) 0.900 (33) 0.904 (32) 0.908 (31) 0.911 (27) 0.924 (24) 1.000 (0)
TREC COVID
Ratio of Manual Annotation Budget ($B/|D|$)
Method $\frac{1}{512}$ $\frac{1}{256}$ $\frac{1}{128}$ $\frac{1}{64}$ $\frac{1}{32}$ $\frac{1}{16}$ $\frac{1}{8}$ $\frac{1}{4}$ $\frac{1}{2}$ all
Depth-k 0.636 (62) 0.673 (66) 0.704 (54) 0.771 (41) 0.818 (29) 0.860 (26) 0.911 (16) 0.953 (14) 0.963 (12) 1.000 (0)
MTF 0.651 (65) 0.614 (40) 0.681 (30) 0.749 (35) 0.833 (28) 0.870 (25) 0.908 (25) 0.949 (16) 0.963 (12) 1.000 (0)
MM-NS 0.572 (81) 0.355 (68) 0.734 (41) 0.806 (33) 0.853 (28) 0.878 (25) 0.873 (25) 0.946 (16) 0.963 (12) 1.000 (0)
CAL(human) 0.452 (80) 0.655 (27) 0.657 (31) 0.860 (28) 0.875 (42) 0.921 (11) 0.937 (13) 0.954 (8) 0.981 (4) 1.000 (0)
CAL(hybrid) 0.751 (43) 0.693 (71) 0.794 (44) 0.663 (35) 0.760 (42) 0.824 (23) 0.901 (29) 0.898 (8) 0.978 (4) 1.000 (0)
SAL(human) 0.456 (56) 0.541 (54) 0.664 (79) 0.348 (74) 0.498 (59) 0.781 (30) 0.817 (20) 0.870 (16) 0.947 (10) 1.000 (0)
SAL(hybrid) 0.285 (82) 0.373 (73) 0.406 (50) 0.742 (47) 0.555 (22) 0.765 (24) 0.806 (19) 0.882 (17) 0.972 (7) 1.000 (0)
LLM-Only 0.950 (19) 0.946 (19) 0.950 (19) 0.950 (19) 0.950 (19) 0.950 (19) 0.950 (19) 0.950 (19) 0.950 (19) 0.950 (19)
Random 0.949 (19) 0.946 (19) 0.950 (19) 0.951 (18) 0.949 (19) 0.952 (18) 0.954 (15) 0.959 (15) 0.977 (12) 1.000 (0)
Naive 0.942 (19) 0.942 (19) 0.944 (19) 0.947 (18) 0.949 (18) 0.957 (18) 0.961 (16) 0.965 (16) 0.972 (15) 1.000 (0)
LARA($n=1$) 0.951 (19) 0.950 (18) 0.951 (18) 0.952 (18) 0.955 (18) 0.956 (17) 0.954 (15) 0.970 (15) 0.986 (11) 1.000 (0)
LARA($n=3$) 0.951 (19) 0.950 (18) 0.951 (18) 0.952 (18) 0.955 (18) 0.956 (17) 0.954 (15) 0.970 (15) 0.986 (11) 1.000 (0)
LARA($n=N$) 0.950 (19) 0.951 (18) 0.952 (18) 0.953 (18) 0.955 (20) 0.955 (15) 0.965 (15) 0.977 (15) 0.990 (11) 1.000 (0)
LLM-Only(rationale) 0.913 (22) 0.913 (22) 0.913 (22) 0.913 (22) 0.913 (22) 0.913 (22) 0.913 (22) 0.913 (22) 0.913 (22) 0.913 (22)
Random(rationale) 0.913 (22) 0.913 (22) 0.914 (22) 0.915 (21) 0.914 (22) 0.916 (21) 0.922 (20) 0.935 (15) 0.955 (14) 1.000 (0)
Naive(rationale) 0.913 (22) 0.914 (22) 0.916 (21) 0.917 (21) 0.919 (21) 0.924 (20) 0.929 (20) 0.937 (16) 0.954 (15) 1.000 (0)
LARA(rationale) 0.915 (22) 0.915 (22) 0.917 (22) 0.918 (21) 0.921 (21) 0.925 (21) 0.930 (18) 0.943 (15) 0.961 (15) 1.000 (0)
LLM-Only(utility) 0.929 (15) 0.929 (15) 0.929 (15) 0.929 (15) 0.929 (15) 0.929 (15) 0.929 (15) 0.929 (15) 0.929 (15) 0.929 (15)
Random(utility) 0.928 (15) 0.930 (15) 0.929 (16) 0.931 (15) 0.930 (15) 0.933 (15) 0.937 (15) 0.950 (15) 0.967 (12) 1.000 (0)
Naive(utility) 0.929 (15) 0.930 (15) 0.929 (16) 0.930 (16) 0.931 (15) 0.935 (15) 0.937 (15) 0.951 (15) 0.971 (12) 1.000 (0)
LARA(utility) 0.930 (15) 0.931 (15) 0.932 (15) 0.932 (15) 0.936 (15) 0.940 (15) 0.947 (15) 0.957 (15) 0.973 (11) 1.000 (0)
LLM-Only (Llama-8B) 0.869 (31) 0.869 (31) 0.869 (31) 0.869 (31) 0.869 (31) 0.869 (31) 0.869 (31) 0.869 (31) 0.869 (31) 0.869 (31)
Random (Llama-8B) 0.869 (31) 0.871 (31) 0.870 (31) 0.872 (31) 0.872 (31) 0.876 (31) 0.883 (31) 0.897 (28) 0.922 (20) 1.000 (0)
Naive (Llama-8B) 0.870 (31) 0.870 (31) 0.871 (31) 0.874 (31) 0.876 (31) 0.881 (31) 0.890 (31) 0.905 (29) 0.946 (16) 1.000 (0)
LARA (Llama-8B) 0.870 (31) 0.871 (31) 0.872 (31) 0.875 (31) 0.876 (31) 0.883 (31) 0.895 (30) 0.917 (27) 0.956 (15) 1.000 (0)
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 2. Comparison of overlap scores of LLM-based methods on different budgets on TREC-7 Ad Hoc (top left), TREC-8 Ad Hoc (top right), TREC Robust04 (bottom left), and TREC-COVID (bottom right).
5. Experimental Results
5.1. Main Results
The main results of the binary relevance experiments on the TREC-7 and TREC-8 Ad Hoc datasets are summarized in Table 2, and the main results of the graded relevance experiments on the TREC Robust04 and TREC COVID datasets are summarized in Table 3. We separate the results to distinguish the binary relevance experiments evaluated using the MAP scores clearly from the graded relevance experiments evaluated using the NDCG scores.
From Tables 2 and 3, we observe that our LARA method outperforms all alternatives on almost any budget constraints on all datasets. We analyze more by answering the following questions.
How does LARA perform compared to existing methods based on manual annotations? Compared to the baseline methods that use only manual judgments (i.e., Depth-k Pooling, MTF, MM-NS, CAL(human), and SAL(human)), LARA almost always performs better, particularly when the budget is low. This result was anticipated because while human annotations are always the ground truth, human-based methods lack the annotation data to make the evaluation stable on a low budget. The baseline hybrid methods (i.e., CAL(hybrid) and SAL(hybrid)) have similar results as those that use only human labels since the regression model can be trained only on small data when the budget is low. Although the Depth-$k$ pooling method does outperform some LARA methods in a few of the scenarios in the binary relevance experiments, we observe that its performance is not consistent, particularly on lower budgets. Moreover, the difference in performance between LARA and the manual-based methods is expected to be larger in real-life applications, since LLMs would be able to assess much more than the limit in these test collections. For example, TREC-8 Ad Hoc is a test collection with depth-100 pooling, but if we use LLMs, we could collect documents from depth-1000 or even from the entire document collection. Indeed, we observe that the LLM-based methods dominate the manual methods on the large TREC Robust04 test collection in Table 3, especially on a lower budget.
How does LARA perform compared to the case of only using LLM? Compared to the case with only LLM annotations (i.e., LLM-Only), LARA’s performance is only marginally better at first, but the difference widens as the budget increases. This is because LARA chooses effective data points to give ground truth judgments, which decreases the number of annotation errors.
How does LARA perform compared to other methods that use both LLM and manual assessments? The Naive and Random methods, which use both LLM and human assessments, almost constantly perform better than the other baseline methods, particularly on a low budget. This alone shows the merit of using both LLM and human judgments. Nonetheless, our LARA method outperforms the Naive and Random methods in most settings, which confirms that our method of calibrating the predictions enhances the use of LLMs for building test collections even more. We elaborate further on this in the exploratory experiments in Section 5.2.
How does the number of annotators affect the performance of LARA? The LARA method with different numbers of annotators (i.e., LARA($n=1$), LARA($n=3$), and LARA($n=N$)), perhaps surprisingly, have almost identical performance on all datasets and budgets, with LARA($n=N$) slightly outperforming the other two in many scenarios. We conjecture that this occurs because as the number of annotators $n$ increases, the number of ground truth annotations would be distributed fairly among each group of topics. This allows each topic to have similar qualities of evaluation results, which helped increase the overall performance. In contrast, when the number of annotators $n$ is small, although the algorithm would be able to choose the most likely-uncertain data point from a wider range of topics, the evaluation performance could vary greatly from topic to topic, since some topics could have very accurate evaluations while others could have poor performance due to differences in manual annotation distribution. This interesting result supports the practicality of our LARA method. We would generally recommend using LARA($n=N$) since it can handle any number of assessors.
How does LARA perform in terms of Maximum Drop? In the tables of main results, we also report the Maximum Drop values in the parentheses. We observe that LARA constantly has one of the lowest drops for all budgets in all datasets. Specifically, compared to the manual methods, LARA’s Maximum Drop is much lower under lower budgets and performs mostly better also under higher budgets. Importantly, only using LLM judgments may lead to very high drops although achieving fairly high Kendall’s $\tau$ scores; this trend is particularly visible in the TREC-8 Ad Hoc experiment with a Maximum Drop of 99. Thus, we argue that it can be risky to rely only on manual judgments on low budgets or to rely only on LLM judgments. We observe that LARA is almost always safer than any alternatives on all budget constraints.
How does the prompt affect the quality of LARA? As an ablation study, we also show results with the rationale prompt (Upadhyay et al., 2024b) and the utility prompt (Thomas et al., 2024). From the results, we observe that LARA outperforms the alternative LLM-based methods on the same prompt and the manual baselines on almost all tested settings. This implies the robustness and consistency of our LARA algorithm.
How does the performance of the LLM affect the quality of LARA? We also show the results of LLM-based methods using Llama-3.1-8B-Instruct, a lightweight variant of Llama-3.1-70B-Instruct. Comparing the LLM-Only methods, we first know that the 8B model performs constantly worse than the 70B model. Nonetheless, while LARA (Llama-8B) generally never outperforms LARA with the 70B models, LARA (Llama-8B) outperforms the rest of the 8B-based models, again showing the effectiveness of the algorithm. Compared to the existing human-based methods, LARA (Llama-8B) performs better under low-budget constraints for all datasets. Thus, when building a dataset under a low budget, LARA is a good option even with a small LLM, but we definitely recommend using a larger and better LLM if available.
5.2. Exploratory Results: LLM Annotation Error
One crucial merit of our algorithm, other than selecting effective documents to manually annotate, is that it can also minimize LLM annotation errors. To visualize this, we compare LARA($n=1$) with Naive and Random methods and show the difference in annotation errors. We do not show LARA($n=3$) and LARA($n=N$) here, because the results were almost identical to LARA($n=1$), so the lines were completely on top of each other. We show the overlap scores of only the LLM-annotated data points. The manually-annotated data points were not considered here, since these data points are always correct and thus not informative.
From the results in Figure 2, we can observe that LARA consistently has a better overlap score than the alternatives on all budgets, which shows that LARA is able to minimize the LLM annotation error by calibrating the LLM judgments with the learned model and by giving truth labels to the data points with the highest expected error. The Naive method also improves its performance because this method also tries to manually annotate the most uncertain data points. Nonetheless, LARA constantly outperforms it in terms of LLM annotation accuracy since the predictions are based on the calibration model $\hat{p}(y|\pi)$ instead of naively using $\pi$.
6. Discussions and Future Directions
As a potentially effective alternative to LARA, one might consider actively fine-tuning the LLM after each batch of human annotations to potentially improve its predictions. However, this approach poses practical challenges. Repeatedly fine-tuning an LLM each time a batch of annotations is collected would be time-consuming, requiring assessors to wait for the model to update and to re-annotate all the data points before proceeding to the next batch [^3] LLM annotation of TREC Robust04 took about 5-6 days; others took about 2 days.. This would greatly disrupt the annotation workflow and reduce efficiency. Instead, our current approach only requires the LLM to output once and focuses on a calibration model that can be updated rapidly [^4] In the experiments, LARA took only about 1-10 seconds to build a full test collection on any budget constraints. Note, however, that we also need to consider the time that takes to assess manually in practice..
However, it may be beneficial to fine-tune the LLM pre-deployment, using historical or domain-specific data. Such pre-deployment tuning could yield a more accurate starting point for LARA, as we see from the results that the quality of LLMs affects the performance of LARA. Exploring the trade-offs between continuous retraining, pre-deployment fine-tuning, and on-the-fly calibration is an interesting avenue for future research. As future work, it would also be valuable to investigate a LARA-like hybrid method on other annotation tasks like nugget matching (Takehi et al., 2023; Nenkova and Passonneau, 2004), which aims to annotate whether some information nugget can be obtained from a text, and e-Discovery (Oard et al., 2018; Cormack and Grossman, 2014), which aims to accurately collect relevant documents for use in legal cases.
7. Conclusions
This paper investigates an effective and efficient way of building a test collection, particularly using both LLM and limited manual assessments. Relying only on limited manual assessments leads to unstable evaluations due to limited collection size, and relying only on LLM judgments is risky and unreliable. Thus, we propose LLM-Assisted Relevance Assessments (LARA) to build a risk-minimized yet large test collection. LARA actively learns to calibrate LLM relevance predictions by strategically selecting documents for human annotation. This learned calibration model guides which documents should be manually reviewed and corrects the LLM predictions for the remaining documents. Experiments on various datasets demonstrate that LARA consistently outperforms all the tested alternatives under almost any budget constraint, suggesting that LARA could help venues like TREC, CLEF, NTCIR, and INEX obtain large, stable, and reliable test collections even when the budget for manual annotations is limited.
Appendix A Prompt
We used the following base prompt for graded relevance, where {description} and {narrative} are given by the NIST assessors.
LLM Prompt Consider the following web page content:
-BEGIN WEB PAGE CONTENT-
{document} -END WEB PAGE CONTENT- A person has typed “{topic}” into a search engine. This person’s intent of this query was “{description}”. Instruction:
Answer if the web content is relevant to the query.
{narrative}. Answer {2} for highly relevant, {1} for somewhat or partially relevant, or {0} for not relevant. Your answer:{
For binary relevance assessments, the final two lines are:
Answer yes or no. Your answer:
Appendix B Proof of Theorem 1
Proof.
Let $p(y=j|\pi)$ be the true relevance probability of relevance level $j$. The Bayes-optimal classifier would predict :
$k=\operatorname*{arg\,max}_{j}p(y=j|\pi).$
Its error is:
$\text{Error}(\pi)=1-p(y=k|\pi).$
Define:
$s=\operatorname*{arg\,max}_{j\neq k}p(y=j|\pi),\quad m=p(y=k|\pi)-p(y=s|\pi)\geq 0.$
Maximizing $\text{Error}(\pi)$ is equivalent to minimizing $p(y=k|\pi)$ under the constraint $p(y=k|\pi)\geq p(y=s|\pi)$. This minimum is achieved at $m=0$, i.e.,
$p(y=k|\pi)=p(y=s|\pi).$
At this point, the classifier is most uncertain. Thus, minimizing $m$ yields the greatest potential reduction in classification error. ∎